ghostmasterTy:

Divide these fractions. Remember to ‘flip’ the second fraction and then multiply. Simplify your answer, if necessary. DO NOT leave your answer as an improper fraction! 10. 7 1/3 ÷ 5 1/5 = 11. 9 6/10 ÷ 4 3/9 = 12. 4 1/12 ÷ 6 1/9 = Add these fractions. First strategy: find a common denominator. You can do this by looking at the smallest denominator and seeing if it will go into the larger denominator evenly. Second strategy: multiply the two denominators together. 13. 3 1/9 x 2 6/9 = (9 should be your Common denominator) 3 1/9 6 6/9

ghostmasterTy:

3 more 25 and im going to be at 100

crispyrat:

ok so 7 1/3 ÷ 5 1/5=22/3 ÷ 26/5 = 22/3 x 5/26 9 6/10 ÷ 4 3/9=96/10 ÷ 39/9=96/10 x 9/39 4 1/12 ÷ 6 1/9=49/12 ÷ 55/9=49/12 x 55/9 Then you just convert those to mixed numbers

ghostmasterTy:

thank you

ghostmasterTy:

but ca you teach me this stuff how to do it

crispyrat:

@crispyrat wrote:
ok so 7 1/3 ÷ 5 1/5=22/3 ÷ 26/5 = 22/3 x 5/26 9 6/10 ÷ 4 3/9=96/10 ÷ 39/9=96/10 x 9/39 4 1/12 ÷ 6 1/9=49/12 ÷ 55/9=49/12 x 55/9 Then you just convert those to mixed numbers
I just converted the numbers to improper and then flipped the second one upside down

ghostmasterTy:

ok

ghostmasterTy:

thank you

ghostmasterTy:

@crispyrat wrote:
@crispyrat wrote:
ok so 7 1/3 ÷ 5 1/5=22/3 ÷ 26/5 = 22/3 x 5/26 9 6/10 ÷ 4 3/9=96/10 ÷ 39/9=96/10 x 9/39 4 1/12 ÷ 6 1/9=49/12 ÷ 55/9=49/12 x 55/9 Then you just convert those to mixed numbers
I just converted the numbers to improper and then flipped the second one upside down
hi